Live From...the Stratosphere
Daily Journal

(From the LFS Archives)

Mission Update:
Flight Day, May 7, 1994

by Juan Rivera, Telescope Operator

Off the ground at 0655Z (2055 Local Time). On board tonight we have a total of 13 people:

We're going to be held to 27,000 for a while by Air Traffic Control for some reason. The flight crew will have to deal with that when the time comes.

0709Z We're passing through 26,000 feet already. I've shut off the flow of liquid nitrogen which is used to pre-cool the telescope cavity. We do this for several reasons: First of all the nitrogen boils off to a very cold dry gas which di splaces the moist air in the cavity. We want the atmosphere in there to be free of any water vapor so it won't fog up the optics and freeze there. Also, we cool the cavity with a huge portable air conditioner and the liquid nitrogen so that it will be col d when we open the aperture door and expose the optics to the ambient temperature at altitude. The mirror started out... Oops... Got'a run!!!!

0749Z Hmm... We'll I spent the last 20 minutes or so attempting to repair a problem with our oscillating secondary mirror. It has been very unreliable lately. Tonight when I turned it on, it blew the main fuse in one of the power supplies that power it. All I could really do in flight was to re-seat all the circuit boards and hope that the problem was being caused by an intermittent connection. Last time this happened I was a hero because I was able to save the mission. This time I was no t so lucky. We are now headed back home and the mission has been aborted. We'll try to get hold of the day crew technicians and see if they can spend some time tonight working on it. Most likely it will be tomorrow before anything can be done. Last time t his happened I found two wires that had been pulled out of the back of a connector on the rear of the chassis. The only problem was that neither wire had anything to do with the problem! It was like trying to find out why your car wouldn't start and findi ng a loose wire that went to the tail lights. It's nice that I found it, but... Anyway, by the time the two broken wires were fixed the problem had mysteriously gone away. It's very very difficult to fix a problem that won't stay bad. We call those "inter mittents". Maybe this time the OSM will stay bad. We call that "inoperative", or "inop" for short. Did you get all that? There will be a test later in the week!

0809Z Time to bag this and secure all the loose equipment and prepare for landing. More next time...

Juan Rivera, Telescope Operator

Back to Journal Index

| LFS Home | Give Us Feedback! | LFS Overview | Search Passport to Knowledge |Passport Home |